# Integer Programming for Generalized Causal Bootstrap Designs

### v2

@@ 3.3 Quadratic-in-treatment estimators
Repeating a similar approach to Section 3.1 for estimators of the form $\hat{\tau}=\sum_{i}b_{i}+\sum_{j}Z_{i}Z_{j}a_{ij}$, we obtain the following quadratic form in the coefficients $a_{ij}$ when computing their variance: $\mathbf{Var}_{Z}[\hat{\tau}]=\sum_{i,j,k,l}a_{ij}a_{kl}\mathbf{Cov}\left[Z_{i}Z_{j},Z_{k}Z_{l}\right]$ We can then repeat the approach of Section 3.1 as long as the coefficients $a_{ij}$ continue to be constant with respect to the treatment assignment and linear in the potential outcomes. In particular, we show in Appendix 7.6 that a linear regression fitting observed outcome $Y^{obs}_{i}$ to the treatment variable $Z$ and a scalar covariate $X$ can be written in this form, with coefficients $a_{ij}$ that are linear in the potential outcomes. One advantage of this approach is that we are no longer required to do sample-splitting or use solely pre-period information to incorporate covariate information and reduce the size of our confidence intervals.
@@ 4.2 Asymptotic Validity
Corollary 1.
For any probabilistic, unconfounded, and individualistic assignment mechanism, Theorem 1 holds with $\beta=8\exp\Big{(}-\frac{\epsilon^{2}}{4}N\tilde{P}\Big{)}+8\exp\Big{(}-\frac{1}{2}N\tilde{P}^{2}\Big{)}\,.$
What of confounded assignments? In fact, we can also prove our approach is asymptotically valid if the assignment is individualistic with bounded confoundedness, as defined by Tan, (2006); De Bartolomeis et al., (2024). Adopting similar notation to Theorem 1,
Theorem 2.
@@
for some $\delta>0$, such that $\delta$ controls the effect of confoundedness, with $\delta=0$ corresponding to unconfoundedness. If the slackness satisfies $\epsilon\geq\delta$, then $\mathbb{P}\left(V^{*}\geq\mathbf{Var}_{Z}[\hat{\tau}]\right)\geq 1-\beta$, where $\beta:=1-8e^{-2N\epsilon_{0}^{2}}-8e^{-N\tilde{P}\epsilon_{1}^{2}/3}$, $\epsilon_{0}=\frac{(\epsilon-\delta)\tilde{P}(1+\delta)}{2+\delta+\epsilon}$, and $\epsilon_{1}=\frac{\epsilon-\delta}{2+\delta+\epsilon}$
Details can be found in Appendix 7.5. We leave the generalization to general confounded assignments to future work.
@@ 5 Simulations on real data
C. R. M. Pairs C. R. M. Pairs C. R. M. Pairs
@@
Table 1: “True CI” and “Median CI” correspond to $95\%$ CI widths in an A/A test under two designs and two estimators. C.R. = Completely Randomized Assignment. M. Pairs = Matched Pairs.
@@
C. R. M. Pairs C. R. M. Pairs C. R. M. Pairs
@@
Table 2: “True CI” and “Median CI” correspond to $95\%$ CI widths for multiplicative effects under two designs and two estimators. C.R. = Completely Randomized Assignment. M. Pairs = Matched Pairs.
@@
To work with a realistic dataset, we look to the International Monetary Fund’s publicly-available Gross Domestic Product (GDP) country-level report for the years 2017–2019 (IMF,, 2024). While a coordinated worldwide randomized experiment affecting GDP is implausible, we can make the more reasonable assumption that some global company’s metric-of-interest in each country is proportional to that country’s GDP, such that the simulation results would approximately hold if such a company were to run a country-level experiment. Our outcome of interest is GDP in 2019, with 2017–2018 GDP acting as pre-period covariates.
@@
Finally, we compare our method to three baselines. Difference-in-means (and its doubly-robust variant) admits conservative, analytical bounds for its sampling variance under both complete randomization and matched-pairs; see Appendix 7.7. We form symmetric, Neyman-style confidence intervals based on these bounds. We next consider the standard sampling bootstrap, which samples with replacement $N$ units $1,000$ times, and reports the empirical confidence intervals as quantiles of the resulting estimator distribution. We also compare to the causal bootstrap method suggested by Aronow et al., (2014); Imbens and Menzel, (2021), which resamples from the isotone copula. This method is only theoretically justified in the complete randomization setting for the difference-in-means estimator. All reported confidence intervals have $95\%$ nominal coverage. We also report the “true confidence intervals” as defined by the $2.5\%$ and $97.5\%$ quantiles of the distribution of the estimate when resampling the assignment according to each design $500$ times.
Scalability
We focus on the top $50$ countries by GDP for our simulation analysis. For a dataset of this size, our integer program took 671 seconds on average for the matched pairs design, and 1057 seconds on average for complete randomization. Matched pairs is faster because of the sparsity of its covariance matrix. However, complete randomization also admits a linear programming formulation for the worst-case copula, which took only 22 seconds on average to solve. We used this faster implementation in our experiments for the sake of expedience. Further details on scalability are provided in Appendix 7.8.
@@
Select results are summarized in Tables 12. A more complete set of results can be found in Appendix 7.1, including implementation details and a reference to the open-source solver. For any experiment with complete randomization, our optimal causal bootstrap mechanism and the isotone copula bootstrap yield the same results down to each simulation as expected: our optimal causal bootstrap recovers the well-known variance-maximizing joint distribution exactly. Across all three inference methods and for the ground truth confidence intervals (CIs), the CIs are greatly reduced by incorporating pre-period information. These CIs are further reduced by the matched-pairs design, with more discussion in Appendix 7.1
All three bootstrap methods achieve lower-than-nominal coverage for the completely randomized design, due to the heavy-tailed nature of GDP data, which causes $\hat{F}_{a}$ to converge slowly. This problem is particularly acute for the completely randomized design, which has high likelihood of assigning the largest GDP countries to the same treatment group. By contrast, the matched-pairs design balances the treated and controlled outcomes, improving the convergence of $\hat{F}_{a}$ and thus the coverage of the sampling bootstrap and our method. We provide more discussion of the observed undercoverage in Section 7.11.

### v3

@@ 3.3 Quadratic-in-treatment estimators
Repeating a similar approach to Section 3.1 for estimators of the form $\hat{\tau}=\sum_{i}b_{i}+\sum_{j}Z_{i}Z_{j}a_{ij}$, we obtain the following quadratic form in the coefficients $a_{ij}$ when computing their variance: $\mathbf{Var}_{Z}[\hat{\tau}]=\sum_{i,j,k,l}a_{ij}a_{kl}\mathbf{Cov}\left[Z_{i}Z_{j},Z_{k}Z_{l}\right]$ We can then repeat the approach of Section 3.1 as long as the coefficients $a_{ij}$ continue to be constant with respect to the treatment assignment and linear in the potential outcomes. In particular, we show in Appendix A.6 that a linear regression fitting observed outcome $Y^{obs}_{i}$ to the treatment variable $Z$ and a scalar covariate $X$ can be written in this form, with coefficients $a_{ij}$ that are linear in the potential outcomes. One advantage of this approach is that we are no longer required to do sample-splitting or use solely pre-period information to incorporate covariate information and reduce the size of our confidence intervals.
@@ 4.2 Asymptotic Validity
Corollary 4.2.
For any probabilistic, unconfounded, and individualistic assignment mechanism, Theorem 4.1 holds with $\beta=8\exp\Big{(}-\frac{\epsilon^{2}}{4}N\tilde{P}\Big{)}+8\exp\Big{(}-\frac{1}{2}N\tilde{P}^{2}\Big{)}\,.$
What of confounded assignments? In fact, we can also prove our approach is asymptotically valid if the assignment is individualistic with bounded confoundedness, as defined by (Tan, 2006; De Bartolomeis et al., 2024). Adopting similar notation to Theorem 4.1,
Theorem 4.3.
@@
for some $\delta>0$, such that $\delta$ controls the effect of confoundedness, with $\delta=0$ corresponding to unconfoundedness. If the slackness satisfies $\epsilon\geq\delta$, then
$\mathbb{P}\left(V^{*}\geq\mathbf{Var}_{Z}[\hat{\tau}]\right)\geq 1-\beta\,,$
where $\beta:=1-8e^{-2N\epsilon_{0}^{2}}-8e^{-N\tilde{P}\epsilon_{1}^{2}/3}$, $\epsilon_{0}=\frac{(\epsilon-\delta)\tilde{P}(1+\delta)}{2+\delta+\epsilon}$, and $\epsilon_{1}=\frac{\epsilon-\delta}{2+\delta+\epsilon}$
Details can be found in Appendix A.5. We leave the generalization to general confounded assignments to future work.
@@ 5 Simulations on real data
Compl. R. Matched Pairs Compl. R. Matched Pairs Compl. R. Matched Pairs
@@
Table 1: “True CI” and “Median CI” correspond to $95\%$ CI widths in an A/A test under two designs and two estimators.
@@
Compl. R. Matched Pairs Compl. R. Matched Pairs Compl. R. Matched Pairs
@@
Table 2: “True CI” and “Median CI” correspond to $95\%$ CI widths for multiplicative effects under two designs and two estimators.
5.1 Empirical Results
@@
To work with a realistic dataset, we look to the International Monetary Fund’s publicly-available Gross Domestic Product (GDP) country-level report for the years 2017–2019, restricted to the top 50 countries by GDP (IMF, 2024). While a coordinated worldwide randomized experiment affecting GDP is implausible, we can make the more reasonable assumption that some global company’s metric-of-interest in each country is proportional to that country’s GDP, such that the simulation results would approximately hold if such a company were to run a country-level experiment. Our outcome of interest is GDP in 2019, with 2017–2018 GDP acting as pre-period covariates.
@@
Finally, we compare our method to three baselines. Difference-in-means (and its doubly-robust variant) admits conservative, analytical bounds for its sampling variance under both complete randomization and matched-pairs; see Appendix A.7. We form symmetric, Neyman-style confidence intervals based on these bounds. We next consider the standard sampling bootstrap, which samples with replacement $N$ units $1,000$ times, and reports the empirical confidence intervals as quantiles of the resulting estimator distribution. We also compare to the causal bootstrap method suggested by (Aronow et al., 2014; Imbens & Menzel, 2021), which resamples from the isotone copula. This method is only theoretically justified in the complete randomization setting for the difference-in-means estimator. All reported confidence intervals have $95\%$ nominal coverage. We also report the “true confidence intervals” as defined by the $2.5\%$ and $97.5\%$ quantiles of the distribution of the estimate when resampling the assignment according to each design $500$ times.
@@
Select results are summarized in Tables 12. A more complete set of results can be found in Appendix A.1, including implementation details and a reference to the open-source solver. For any experiment with complete randomization, our optimal causal bootstrap mechanism and the isotone copula bootstrap yield the same results down to each simulation as expected: our optimal causal bootstrap recovers the well-known variance-maximizing joint distribution exactly. Across all three inference methods and for the ground truth confidence intervals (CIs), the CIs are greatly reduced by incorporating pre-period information. These CIs are further reduced by the matched-pairs design, with more discussion in Appendix A.1
All three bootstrap methods achieve lower-than-nominal coverage for the completely randomized design, due to the heavy-tailed nature of GDP data, which causes $\hat{F}_{a}$ to converge slowly. This problem is particularly acute for the completely randomized design, which has high likelihood of assigning the largest GDP countries to the same treatment group. By contrast, the matched-pairs design balances the treated and controlled outcomes, improving the convergence of $\hat{F}_{a}$ and thus the coverage of the sampling bootstrap and our method. We provide more discussion of the observed undercoverage in Section A.10.
@@
5.2 Scalability Analysis
Figure 1: Runtime scaling for matched pairs, integer program formulation.
Figure 2: Runtime scaling for complete randomization, linear program formulation.
As mentioned, we focus on the top $50$ countries by GDP for our simulation analysis. For a dataset of this size, our integer program took 671 seconds on average for the matched pairs design, and 1057 seconds on average for complete randomization. Matched pairs is faster because of the sparsity of its covariance matrix. However, complete randomization also admits a linear programming formulation for the worst-case copula, which took only 22 seconds on average to solve. We used this faster implementation in our experiments for the sake of expedience.
To assess the scalability of our formulation, we solved for the worst-case copula under a null effect, varying the number of units. Figures 22 presents the results of our scaling analysis, where each data point in the plot represents an average over $10$ simulation runs, along with $95\%$ confidence intervals based on standard errors.